# Using Uncertainty Quantification to Characterize and Improve Out-of-Domain Learning for PDEs

### v1

@@ F.1.4 Hyperbolic Linear Advection Equation
Figures 31-34 show the solution profiles for the hyperbolic linear advection equation conditions for an in-domain task, and for small, medium and large OOD shifts, respectively, of the velocity parameter $\beta>0$. The solution is a rightward moving shock, where $\beta$ controls the shock speed. We see that in this difficult shock case the baseline methods, e.g., BayesianNO, VarianceNO and MC-DropoutNO suffer from numerical artifacts of artificial oscillations, being over-diffusive and lagging of the shock position. Only EnsembleNO and our DiverseNO accurately capture the shock location and have the largest uncertainty there. See corresponding metric results in Table 6.
Table 6: 1-d Linear Advection MSE normal-$\downarrow$, NLL normal-$\downarrow$, n-MeRCI normal-$\downarrow$, RMSCE normal-$\downarrow$ and CRPS normal-$\downarrow$ (mean and standard deviation over 5 seeds) metrics for different UQ methods on the 1-d linear advection equation in-domain and with small, medium and large OOD shifts, where $\beta^{\text{train}}\in[1,2]$. Bold indicates values within one standard deviation of the best mean.
@@ F.2 Cost performance curves
Figure 35: Cost-performance tradeoff. MSE normal-$\downarrow$ and n-MeRCI normal-$\downarrow$ vs number of parameters for EnsembleNO and DiverseNO with varying number of parameters. for (a-b) 1-d Heat equation, medium OOD shift, i.e., $k^{\text{train}}\in[1,5],k^{\text{test}}\in[6,7]$ and (c-d) 1-d PME, medium OOD shift, i.e., $m^{\text{train}}\in[2,3],m^{\text{test}}\in[4,5]$.
@@ F.3 Effect of ProbConserv update
Table 7: Effect of ProbConserv update, 1-d heat equation. MSE normal-$\downarrow$, Conservation Error (CE) (should be zero) and n-MeRCI normal-$\downarrow$ (mean and standard deviation over 5 seeds) for different UQ methods with and without ProbConserv evaluated on 1-d heat equation with small, medium and large OOD shifts, where $k^{\text{train}}\in[1,5]$. Bold indicates values within one standard deviation of the best mean.
@@
Table 8: Effect of ProbConserv update, 1-d PME. MSE normal-$\downarrow$, Conservation Error (CE) (should be zero) and n-MeRCI normal-$\downarrow$ (mean and standard deviation over 5 seeds) for different UQ methods with and without ProbConserv evaluated on 1-d PME with small, medium and large OOD shifts, where $m^{\text{train}}\in[2,3]$. Bold indicates values within one standard deviation of the best mean.
@@
In this section, we report the solution profiles and metrics, which compare the effect of using the various UQ estimates for NOs within the ProbConserv framework. Tables 7-9 show that the conservation error (CE) after applying ProbConserv is 0, and that the CE for the unconstrained models can be quite large, especially on the harder tasks. The results on the “easy” heat equation task in Figure 36 and corresponding Table 7 show that ProbConserv improves the MSE for all the UQ methods. The results on the “harder” problems in Figure 37 and corresponding Table 8 for the PME and Figure 38 and corresponding Table 9 for the Stefan equation show the impact of good uncertainty estimates as input to ProbConserv. We see that ProbConserv improves the solution profiles and MSE for the methods except for VarianceNO, whose UQ estimate is not well-correlated with the error.
Table 9: Effect of ProbConserv update, 1-d Stefan. MSE normal-$\downarrow$, Conservation Error (CE) (should be zero) and n-MeRCI normal-$\downarrow$ (mean and standard deviation over 5 seeds) for different UQ methods with and without ProbConserv evaluated on the 1-d Stefan equation with small, medium and large OOD shifts, where $u{{}^{*}}^{{\text{train}}}\in[0.6,0.65]$. Bold indicates values within one standard deviation of the best mean.

### v2

@@ F.1.4 Hyperbolic Linear Advection Equation
We show results for the following two tasks: constant parameter input by varying the speed $\beta$ and non-constant initial condition input by varying the initial shock location $a$.
Constant parameter to solution mapping.
Figures 36-39 show the solution profiles for the hyperbolic linear advection equation conditions for an in-domain task, and for small, medium and large OOD shifts, respectively, of the velocity parameter $\beta>0$. The solution is a rightward moving shock, where $\beta$ controls the shock speed. We see that in this difficult shock case the baseline methods, e.g., BayesianNO, VarianceNO and MC-DropoutNO suffer from numerical artifacts of artificial oscillations, being over-diffusive and lagging of the shock position. Only EnsembleNO and our DiverseNO accurately capture the shock location and have the largest uncertainty there. See corresponding metric results in Table 6.
Table 6: 1-d Linear Advection MSE $\downarrow$, NLL $\downarrow$, n-MeRCI $\downarrow$, RMSCE $\downarrow$ and CRPS $\downarrow$ (mean and standard deviation over 5 seeds) metrics for different UQ methods on the 1-d linear advection equation in-domain and with small, medium and large OOD shifts, where $\beta^{\text{train}}\in[1,2]$. Bold indicates values within one standard deviation of the best mean.
@@
Non-constant initial condition to solution mapping.
Here we test non-constant function input to the NO by varying the initial condition through the initial shock location $a$. We see that every method gives good predictions in-domain (Figure 40) and has low in-domain MSE (Table 7). Table 7 also shows that DiverseNO performs best or second-best to EnsembleNO in MSE, while being computationally cheaper. With respect to the n-MeRCI metric, Table 8 shows that DiverseNO performs around 1.2×1.2$\times$1.2 × to 1.5×1.5$\times$1.5 × better than EnsembleNO and 2.7×2.7$\times$2.7 × to 7.4×7.4$\times$7.4 × better than other baselines on OOD inputs.
(a) Input
(b) BayesianNO
(c) MC-DropoutNO
(d) EnsembleNO
(e) DiverseNO
Figure 40: 1-d Linear Advection (non-constant input), in-domain. Uncertainty estimates (3 standard deviations) from various UQ methods in-domain where $a^{\text{train}}\in[0.45,0.55]$.
(a) Input
(b) BayesianNO
(c) MC-DropoutNO
(d) EnsembleNO
(e) DiverseNO
Figure 41: 1-d Linear Advection (non-constant input), small OOD shift. Uncertainty estimates (3 standard deviations) from various UQ methods with small OOD shift with $a^{\text{train}}\in[0.45,0.55]$ and $a^{\text{test}}\in[0.4,0.45]$.
(a) Input
(b) BayesianNO
(c) MC-DropoutNO
(d) EnsembleNO
(e) DiverseNO
Figure 42: 1-d Linear Advection (non-constant input), medium OOD shift. Uncertainty estimates (3 standard deviations) from various UQ methods with medium OOD shift with $a^{\text{train}}\in[0.45,0.55]$ and $a^{\text{test}}\in[0.6,0.65]$.
(a) Input
(b) BayesianNO
(c) MC-DropoutNO
(d) EnsembleNO
(e) DiverseNO
Figure 43: 1-d Linear Advection (non-constant input), large OOD shift. Uncertainty estimates (3 standard deviations) from various UQ methods with large OOD shift with $a^{\text{train}}\in[0.45,0.55]$ and $a^{\text{test}}\in[0.35,0.4]$.
Figures 41, 42 and 43 show the predictions of the various methods under small, medium and large OOD shifts. All methods exhibit high errors near the shock location and at the boundary, particularly for medium (Figure 42) and large OOD shifts (Figure 43). BayesianNO and MC-DropoutNO output very low uncertainty despite these high errors. In contrast, the UQ from DiverseNO is correlated with the error and is highest near the shock and the right boundary.
Table 7: 1-d Linear Advection (non-constant input). MSE $\downarrow$ (mean and standard deviation over 5 seeds) for different UQ methods on the 1-d linear advection equation in-domain and with small, medium and large OOD shifts. Input is a non-constant function $u(x,0)=1_{x\leq a}$ with $a^{\text{train}}\in[0.45,0.55]$. Bold indicates values within one standard deviation of the best mean.
Models In-domain OOD-small OOD-medium OOD-large
$a^{\text{test}}\in[0.45,0.55]$ $a^{\text{test}}\in[0.4,0.45]$ $a^{\text{test}}\in[0.6,0.65]$ $a^{\text{test}}\in[0.35,0.4]$
BayesianNO 1.3e-06 (1.2e-06) 3.3e-04 (6.3e-05) 4.9e-03 (1.4e-03) 2.3e-02 (3.2e-03)
MC-DropoutNO 2.7e-05 (5.2e-06) 4.2e-04 (6.6e-05) 2.9e-03 (5.4e-04) 2.9e-02 (4.9e-03)
EnsembleNO 7.5e-07 (6.5e-07) 3.1e-04 (5.0e-05) 4.3e-03 (9.1e-05) 2.2e-02 (1.5e-03)
DiverseNO 1.5e-06 (1.1e-06) 3.4e-04 (9.6e-05) 2.7e-03 (7.6e-04) 2.8e-02 (5.5e-03)
Table 8: 1-d Linear Advection (non-constant input). n-MeRCI $\downarrow$ (mean and standard deviation over 5 seeds) for different UQ methods on the 1-d linear advection equation in-domain and with small, medium and large OOD shifts. Input is a non-constant function $u(x,0)=1_{x\leq a}$ with $a^{\text{test}}\in[0.45,0.55]$. Bold indicates values within one standard deviation of the best mean.
Models In-domain OOD-small OOD-medium OOD-large
$a^{\text{test}}\in[0.45,0.55]$ $a^{\text{test}}\in[0.4,0.45]$ $a^{\text{test}}\in[0.6,0.65]$ $a^{\text{test}}\in[0.35,0.4]$
BayesianNO 0.75 (0.39) 1.06 (0.01) 0.94 (0.20) 0.87 (0.18)
MC-DropoutNO 0.65 (0.19) 1.00 (0.01) 0.89 (0.18) 0.86 (0.18)
EnsembleNO 0.18 (0.07) 0.19 (0.03) 0.19 (0.05) 0.37 (0.11)
DiverseNO 0.22 (0.15) 0.16 (0.02) 0.12 (0.07) 0.31 (0.10)
F.1.5 Elliptic 2-d Darcy Flow
Here we provide an elliptic, steady-state 2-d test case, i.e., Darcy Flow, where the solution $u(x)$ denotes the unknown pressure and $k$ the constant permeability field. Table 9 shows the MSE metric for all methods in-domain and across different OOD shifts. The in-domain MSE is $\approx 10^{-11}$ for all methods and increases by $10^{4}$ for the largest OOD shift. (Table 9) also shows that the OOD MSE of DiverseNO is improved (1.2$\approx$ 1.2$\times$ 1.2 ×) upon that of the competing baselines. With respect to the meaningful n-MeRCI metric that measures error correlation with uncertainty estimates, Table 10 shows that EnsembleNO and DiverseNO perform 10$\approx$ 10$\times$ 10 × better than the other methods with DiverseNO being more computationally efficient. Similar trends hold for the CRPS metric in Table 11, where DiverseNO and EnsembleNO outperform other baselines by 1.5×1.5$\times$1.5 ×. These 2-d results are consistent with 1-d experiments.
Figure 44 illustrates the solution profile (left column), absolute error (middle column) and uncertainty plots (right column) under a large OOD shift for our DiverseNO model and the baselines. The absolute errors from all models except VarianceNO are concentrated around the center of the domain where the pressure values in the solution profile are highest. We see that the errors from DiverseNO are slightly lower than that of the baselines. BayesianNO (Figure 44(a)) and VarianceNO (Figure 44(b)) output uncertainty estimates spread uniformly over the domain. Uncertainty estimates from MC-DropoutNO (Figure 44(c)) are highest near the boundary instead of the center region. Uncertainty estimates from DiverseNO (Figure 44(e)) are highest around the center, correlating better with the error.
(a) BayesianNO
(b) VarianceNO
(c) MC-DropoutNO
(d) EnsembleNO
(e) DiverseNO
Figure 44: 2-d Darcy Flow, large OOD shift. Solution profiles (left column), absolute error (middle column) and uncertainty estimates (right column) from different UQ methods under a large OOD shift ($k^{\text{train}}\in[3,4]$ and $k^{\text{test}}\in[5,6]$).
Table 9: 2-d Darcy Flow. MSE $\downarrow$ for different UQ methods on the 2-d Darcy Flow equation in-domain and with small, medium and large OOD shifts, where $k^{\text{train}}\in[3,4]$. Bold indicates the best mean.
Models In-domain OOD-small OOD-medium OOD-large
$k^{\text{test}}\in[3,4]$ $k^{\text{test}}\in[4,4.5]$ $k^{\text{test}}\in[4.5,5]$ $k^{\text{test}}\in[5,6]$
BayesianNO 4.3e-11 4.6e-09 5.6e-08 4.2e-07
VarianceNO 7.7e-12 7.4e-10 2.9e-08 1.0e-06
MC-DropoutNO 1.2e-09 4.7e-09 5.4e-08 4.4e-07
EnsembleNO 3.9e-11 4.2e-09 5.0e-08 3.4e-07
DiverseNO 3.9e-11 3.8e-09 4.4e-08 2.9e-07
Table 10: 2-d Darcy Flow. n-MeRCI $\downarrow$ for different UQ methods on the 2-d Darcy Flow equation in-domain and with small, medium and large OOD shifts, where $k^{\text{train}}\in[3,4]$. Bold indicates the best mean.
Models In-domain OOD-small OOD-medium OOD-large
$k^{\text{test}}\in[3,4]$ $k^{\text{test}}\in[4,4.5]$ $k^{\text{test}}\in[4.5,5]$ $k^{\text{test}}\in[5,6]$
BayesianNO 0.47 0.79 0.83 0.71
VarianceNO 0.28 0.78 0.87 0.78
MC-DropoutNO 0.24 0.86 0.88 0.75
EnsembleNO 0.11 0.03 0.13 0.26
DiverseNO 0.05 0.06 0.07 0.24
Table 11: 2-d Darcy Flow. CRPS $\downarrow$ for different UQ methods on the 2-d Darcy Flow equation in-domain and with small, medium and large OOD shifts, where $k^{\text{train}}\in[3,4]$. Bold indicates the best mean.
Models In-domain OOD-small OOD-medium OOD-large
$k^{\text{test}}\in[3,4]$ $k^{\text{test}}\in[4,4.5]$ $k^{\text{test}}\in[4.5,5]$ $k^{\text{test}}\in[5,6]$
BayesianNO 0.034 1.89 5.17 10.36
VarianceNO 0.430 1.59 8.43 30.70
MC-DropoutNO 0.031 1.70 5.02 10.68
EnsembleNO 0.033 1.27 4.05 7.40
DiverseNO 0.035 1.28 4.01 7.19
@@ F.2 Cost performance curves
Figure 45: Cost-performance tradeoff. MSE $\downarrow$ and n-MeRCI $\downarrow$ vs number of parameters for EnsembleNO and DiverseNO with varying number of parameters. for (a-b) 1-d Heat equation, medium OOD shift, i.e., $k^{\text{train}}\in[1,5],k^{\text{test}}\in[6,7]$ and (c-d) 1-d PME, medium OOD shift, i.e., $m^{\text{train}}\in[2,3],m^{\text{test}}\in[4,5]$.
@@ F.3 Effect of ProbConserv update
In this section, we report the solution profiles and metrics, which compare the effect of using the various UQ estimates for NOs within the ProbConserv framework. Tables 12-14 show that the conservation error (CE) after applying ProbConserv is 0, and that the CE for the unconstrained models can be quite large, especially on the harder tasks. The results on the “easy” heat equation task in Figure 46 and corresponding Table 12 show that ProbConserv improves the MSE for all the UQ methods. The results on the “harder” problems in Figure 47 and corresponding Table 13 for the PME and Figure 48 and corresponding Table 14 for the Stefan equation show the impact of good uncertainty estimates as input to ProbConserv. We see that ProbConserv improves the solution profiles and MSE for the methods except for VarianceNO, whose UQ estimate is not well-correlated with the error.
Table 12: Effect of ProbConserv update, 1-d heat equation. MSE $\downarrow$, Conservation Error (CE) (should be zero) and n-MeRCI $\downarrow$ (mean and standard deviation over 5 seeds) for different UQ methods with and without ProbConserv evaluated on 1-d heat equation with small, medium and large OOD shifts, where $k^{\text{train}}\in[1,5]$. Bold indicates values within one standard deviation of the best mean.
@@
Table 13: Effect of ProbConserv update, 1-d PME. MSE $\downarrow$, Conservation Error (CE) (should be zero) and n-MeRCI $\downarrow$ (mean and standard deviation over 5 seeds) for different UQ methods with and without ProbConserv evaluated on 1-d PME with small, medium and large OOD shifts, where $m^{\text{train}}\in[2,3]$. Bold indicates values within one standard deviation of the best mean.
@@
Table 14: Effect of ProbConserv update, 1-d Stefan. MSE $\downarrow$, Conservation Error (CE) (should be zero) and n-MeRCI $\downarrow$ (mean and standard deviation over 5 seeds) for different UQ methods with and without ProbConserv evaluated on the 1-d Stefan equation with small, medium and large OOD shifts, where $u{{}^{*}}^{{\text{train}}}\in[0.6,0.65]$. Bold indicates values within one standard deviation of the best mean.